OpenStudy (star):

Let T: R^2 -> R^2 be a linear transformation such that T(i) = (3 2) and T(j) = (-6 -4) Find a matrix that induces T Is T invertible?

OpenStudy (star):

T(i) = \[\left(\begin{matrix}3 \\ 2\end{matrix}\right)\] T(j) = \[\left(\begin{matrix}-6 \\ -4\end{matrix}\right)\]

OpenStudy (anonymous):

Not too clear on the notation, I will guess that u want \[\left( \begin{array}{cc} 3 & -6 \\ 2 & -4 \end{array} \right)\] which will map(i,j) to (3i -6j,2i-4j) = i(3,2) +j(-6,-4) Since the determinant is not 0, this transformation is invertible.

OpenStudy (zarkon):

(3*-4)-(2*(-6))=0

OpenStudy (anonymous):

Oops, quite right, so not invertible:-)